OpenStudy (anonymous):

2 cos^2x+2sin x-12=0, interval 0-2TT, i need help :(

OpenStudy (anonymous):

start with \[2(1-\sin^2(x))+2\sin(x)-12=0\] then solve the quadratic equation in \(\sin(x)\)